News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 23 November 2001 2.30pm EST

Last week the observing schedule was halted twice due to high radiation associated with solar flare activity. The loads were halted on Nov 19 at 9:50pm EST through a ground command to execute the SI Safing SCS 107. This ensured that the accumulated radiation dose for ACIS remained below the allowed threshold. Loads were restarted at 9.00am EST on Nov 21. The second interruption occurred on Nov 22 at 8:45pm EST from an autonomous trip of the radiation monitor and execution of SCS 107. In both cases, all spacecraft actions completed nominally. A contingency maneuver load was uplinked at 3:15am EST on Nov 23 following the second event to avoid an Earth occultation during the next perigee. Observations will be resumed once the radiation reaches an acceptable level and the targets impacted by the events will be rescheduled.

Chandra passed through the 2001 Leonid meteor shower on Nov 19 without incident. The aspect system reacquired stars as expected after the 3 hour gyro hold during an Earth occultation near perigee. Two detailed articles about Chandra and the Leonids can be found in the Chandra Chronicles at http://chandra.harvard.edu/chronicle/index.html.

The schedule of observations for the coming week (assuming a restart following the solar stoppage) is shown below and includes a series of grating observations of PKS2155-304 coordinated with XMM-Newton.

------------------------------------------
SMC Field 7            ACIS-I       Nov 25
Vela Pwn               ACIS-S
MACS-J0329.6-0211      ACIS-I
Radiation Belts
PSS0133+0400           ACIS-S       Nov 26
UV Ceti, Gliese 65     HRC-S/LETG
PKS2155-304            ACIS-I/LETG  Nov 27
Vela Pwn               ACIS-S
PKS2155-304            ACIS-I/LETG  Nov 28
V1494 Aql              HRC-S/LETG
Radiation Belts
FXRT011030 (TOO)       ACIS-S       Nov 29
MACSJ0047.3-0810       ACIS-I
PKS2155-304            ACIS-I/LETG
PKS2155-304            HRC-S/LETG
PKS2155-304            ACIS-S/LETG  Nov 30
PKS2155-304            ACIS-S/HETG
PKS2155-304            ACIS-I/LETG  Dec 1 
Radiation Belts
Her X-1 Main HI Late   ACIS-S/HETG
Vela Pwn               ACIS-S       Dec 2

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports